OpenStudy (maddy1251):

Need help with a Log? Will fan and medal. Find y if y= log32 1. Not too sure if y=0, my one calculator shot that out. But I don't know how to quite explain that.

OpenStudy (maddy1251):

@superdavesuper

OpenStudy (maddy1251):

@ganeshie8

OpenStudy (math&ing001):

Here's a hint : \[\log _{32}(x)=\frac{ \ln(x) }{ \ln(32) }\]

OpenStudy (maddy1251):

Are those natural logs? the fraction part.

OpenStudy (math&ing001):

Yes this is normal log, the inverse of exponential, not the log base 10.

OpenStudy (maddy1251):

Okay. Now would this be the same as \[\frac{ \ln(1) }{ \ln(32) } = 0?\]

OpenStudy (math&ing001):

Yeah !